This is NA61/SHINE electronic logbook (read-only BACKUP!)

Run summary Shift logbook Run settings Beam logbook BPD logbook DAQ logbook DCS logbook Experts logbook Gas logbook TOF logbook Trigger logbook PSD logbook VD logbook Run coordination Detector Upgrade DRS logbook
  Logbook for all experts' activities  NA61/SHINE logo
Message ID: 326     Entry time: Thu, 7.07.2022 16:46:37
Author: Y. Nagai 
Subject: Beam position fine tunings (T2K replica target run 2022) 
We observe beam vertical position change frequently. Below summarize the adjustment history.
It seems to have some correlation with accelerator performance, but not clear.

===================================================================================================
Y (reference) = 2.6 cm
Y (before) = observed shifted beam position (vertical)
Y (after) = beam position after re-tuning (vertical)
===================================================================================================
time | Y (before) | Y (after) | TRIM05 change | Comment
===================================================================================================
06.07 22:55 | 2.2 cm | 2.6 cm | +2.0A --> -2.0A | Beam intensity was unstable. YN did.
07.07 02:23 | 2.0 cm | 2.6 cm | -2.0A --> -4.5A | After intensity raised by CCC. YN did.
07.07 16:48 | 2.82 cm | 2.6 cm | -4.5A --> -3.0A | After short accelerator break. YN did.
07.07 18:44 | 2.97 cm | 2.57 cm | -3.0A-> -2.0A-> -1.0A-> 0.0A | Several short accelerator breaks. TN did. -3.0A (2.9cm) -> -2.0A (2.93cm) -> -1.0A (2.70cm) -> 0.0A (2.57cm)
07.07 21:37 | 2.32 cm | 2.6 cm | 0.0A -> -2.0A | Beam intensity is higher (S1 = 48K). It goes up from 40K to 48K around 19:00. Then at 20:00, Y position shift is observed.
14.07 14:30 | 3.0 cm | 2.7 cm | -2.0A -> +1.0A | Beam intensity is about S1 = 43K
19.07 12:00 | 3.1 cm | 2.8 cm | -2.0A -> 0.0A | Beam intensity is S1=61K after accelerator trouble last night. Maybe the TRIM05 starting value was reset to -2.0A after hard reset during access?
19.07 13:00 | 2.8 cm | 2.6 cm | 0.0A -> +1.0A | Fine adjust to hit 2.6 cm position
20.07 19:45 | 2.9 cm | 2.5 cm | +1.0A -> +2.0A | After MD, we observed Y shift.
ELOG V3.1.3-7933898